find the locus of a point whose sum of distances from the origin and the line x-2 is 4 units.

Grade:11

1 Answers

Vijay Mukati
8 years ago
Dear Student,

let the point be (h,k),
Therefore, according to conditions,.
square_root (h2 + k2) + (2-h) = 4,
or
h2 + k2 = (2+h)2
k2 = 4(h+1),
or y2 = 4(x+1), Required locus is the parabola.

Thanks,